# Data Analysis Prediction over Multiple Unseen Datasets: A Vector Embedding Approach

### v2

@@ III-C Dataset Selection
The selection of the most similar datasets has been implemented using three different approaches. Different similarity functions are easily plugged into our pipeline. The first method uses the cosine similarity to determine how similar two datasets are in their vector embedding space $z$. Cosine similarity calculates the angle between the two vectors and a measure of similarity that is independent of the vector magnitudes. The higher its value, the most similar are the two datasets. The following equation describes the cosine similarity between vectors $A$ and $B$:
$\cos\left(\theta\right)=\dfrac{A\cdot B}{\left\|A\right\|_{2}\left\|B\right\|_{2}}=\dfrac{\sum_{i=1}^{n}A_{i}B_{i}}{\sqrt{\sum_{i=1}^{n}A^{2}_{i}}\sqrt{\sum_{i=1}^{n}B^{2}_{i}}}$ (5)
The alternative method involves calculating the Euclidean distance between the two vectors in order to determine how similar the two datasets are. Euclidean distance is particularly useful for capturing the geometric closeness between vectors and finding the most relevant datasets based on the organisation’s requirements. The Euclidean distance between vectors $\mathbf{u},\mathbf{v}$ is defined as:
$d(\mathbf{u},\mathbf{v})=\sqrt{\sum_{i=1}^{n}(u_{i}-v_{i})^{2}}$ (6)
@@
The smaller the distance value the most similar are the datasets. Dataset selection using cosine similarity or euclidean distance selects a percentage of $\lambda,\lambda>0$ of the closest datasets to dataset $D_{o}$. The third approach involves utilising the K-Means [23, 24] clustering technique to choose the most relevant datasets. The datasets from $D$ are divided into $s$ ($s>1$) separate clusters, where datasets with similar features are grouped together in the same cluster based on similarity equations. The silhouette scores [25] are used to determine the number of clusters $s$ for each possible value of $S$. This is done by the following equation:
$SilhouetteScore(v_{i})=\frac{b(v_{i})-a(v_{i})}{\max(a(v_{i}),b(v_{i}))}$ (7)
@@
where for each vector $v_{i}$ finds the mean intra-cluster distance ($a(v_{i})$) which is the distance with the other vectors in the same cluster, and the mean nearest-cluster distance ($b(v_{i})$) is the minimum average distance with the other vectors in the different cluster. Silhouette Score ranges from $-1$ to $1$ and the higher value defines the best $s$ number for clusters.
@@
Algorithm 1, shown how k-means clustering is created and how the most relevant datasets $D_{r}$ are selected based on the target dataset $D_{o}$. Using the most optimal $s$ value with the highest Silhouette score, K-means clustering is performed on the vector representations $V$ of each dataset, grouping datasets with same features into the same cluster. Next, the algorithm uses the vector $v_{o}$ of dataset $D_{o}$ to find the closest cluster centroid. All datasets in that cluster are defined as the relevant datasets $D_{r}$. These datasets are then used to model the analytics operator at hand. However, we defined a maximum and minimum size for $D_{r}$, and if these conditions are not met, datasets are either removed from the cluster or added based on their distance from the cluster centroid. These small adjustments are only made in cases where the clustering technique does not yield results that satisfy our requirements.
IV Evaluation
@@
We compared our framework with Apollo [12, 13], which models analytic operators using data content. Two loss functions to measure prediction accuracy are employed: root-mean-square error (RMSE) and mean absolute error (MAE). RMSE is sensitive to outliers, while MAE is not; conversely, RMSE accounts for error direction, which MAE cannot. Speedup metrics are also used to evaluate how efficiently our framework models operator $\Phi$. Specifically, Speedup and Amortized Speedup measure the time required to approximate each operator versus exhaustively executing them on all datasets. Speedup is defined as $\frac{T{{}^{(i)}_{op}}}{T{{}^{(i)}_{SimOp}+T_{vec}+T_{sim}+T_{pred}}}$, where $T{{}^{(i)}_{op}}$ is the time to execute operator $i$ on all datasets, $T{{}^{(i)}_{SimOp}}$ is the time to model the operator with datasets from similarity search, $T_{vec}$ is the vector embedding computation time, $T_{sim}$ is the similarity search time, and $T_{pred}$ is the prediction time for $D_{o}$. Amortized speedup includes dataset vectorization, performed once per data lake for multiple operators (in our case two operators). We also evaluate our dataset vectorization model, NumTabData2Vec, which projects datasets into vector embedding space $z$. Three versions were built with vector sizes of $100$, $200$, and $300$, each featuring eight transformer layers. The models were trained for 50 epochs on four NVIDIA A10 GPUs in parallel.
IV-A Evaluation Setup
@@
Our framework is deployed over an AWS EC2 virtual machine server running with 48 VCPUs of AMD EPYC 7R32 processors at 2.40GHz, and four A10s GPUs with 24GB of memory each, $192GB$ of RAM memory, and $2TB$ of storage, running over Ubuntu 24.4 LTS. Our code is written in Python (v.3.9.1) and PyTorch modules (v.2.4.0). Apollo was deployed in a virtual machine with 8 VCPUs Intel Xeon E5-2630 @ 2.30GHz, $64GB$ of RAM memory, and $250GB$ of storage, running Ubuntu 24.4 LTS like in their experimental evaluation.
IV-B Datasets
TABLE I: Dataset properties for experimental evaluation
@@ IV-B Datasets
We evaluated our framework using four diverse datasets to represent real-world scenarios. Table I summarizes these datasets and their attributes. The vectorization module, NumTabData2Vec, was trained on data separate from the experimental evaluation data, split $60\%$ for training and $40\%$ for testing. The Household Power Consumption (HPC) dataset [28] contains electric power usage measurements from a household in Sceaux, France. It includes $401$ datasets, each with $2051$ tuples and seven features recorded at one-minute intervals. The Adult dataset [29], commonly used for binary classification, predicts whether an individual earns more or less than $50K$ annually. It comprises $100$ datasets, each with $228$ individuals and various socio-economic features. The Stock Market dataset [30] includes daily NASDAQ stock prices obtained from Yahoo Finance, with $508$ datasets. Each dataset contains $13$ to $1959$ tuples, each describing seven feature attributes. The Weather dataset [31] provides hourly weather measurements from $36$ U.S. cities between $2012$ and $2017$, split into $49$ datasets, each with $516$ tuples and seven features.
(a) Linear Regression RMSE error loss
(b) Linear Regression MAE error loss
(c) MLP for Regression RMSE error loss
(d) MLP for Regression MAE error loss
@@
Our framework was evaluated by registering the accuracy of predicting the output of various ML operators over multiple datasets in $D$ without actually executing the operator on them. To evaluate our scheme and its parameters, we use all four datasets, ranging the size of the produced vectors as well as the similarity functions used. We project all datasets into $k$-dimensional spaces with varying vector dimensions ($100$, $200$, and $300$). For each dataset in Table I, we model different operators: For the regression datasets (Household Power Consumption and Stock Market), we model Linear Regression (LR) and Multi-Layer Perceptron (MLP) operators; for the classification datasets (Weather and Adult), we model the Support Vector Machine (SVM) and MLP classifier operators. Each experiment has been executed $10$ times and we report the average of the error loss, as well as the speedup.
(a) Linear Regression RMSE error loss
(b) Linear Regression MAE error loss
(c) MLP for Regression RMSE error loss
(d) MLP for Regression MAE error loss
@@
IV-C Evaluation Results
Figures 3, 4, 5, and 6 present the evaluation results for each method, comparing the performance of different similarity search techniques across various vector embedding representation spaces. The red (with hatches), brown, and blue bars correspond to vector embeddings of size 100, 200, and 300 respectively. In each sub-figure, the y-axis represents the error loss value, while the x-axis displays the similarity search method applied over the vector embeddings. Figures 3 and 4 show the results for the Stock market and Household power consumption datasets, where the bottom sub-figure demonstrates the MLP regression model, and the top sub-figure presents the LR model. Figures 5 and 6 depict the evaluation results for the Weather and Adult datasets. In these Figures, the top sub-figure shows the SVM with SGD results, while the bottom sub-figure shows the MLP classifier. The left sub-figures in all Figures use the RMSE loss function, whereas the right sub-figures use the MAE loss function.
Figure 3, we show, for the HPC dataset, shows as increase the vector dimension size there is slightly lower prediction error for all the operator modelling. While for different similarity methods did not result in any significant differences in the prediction error loss for all the operator modelling. This suggests that, regardless the similarity selection method, our framework effectively selects the most optimal subset of data to improve model predictions on the unseen input dataset $D_{o}$. Additionally, we observe higher error loss with a vector size of 100, which can be attributed to the reduced representation capacity of lower-dimensional vectors. This limitation results in fewer “right” datasets being selected.
@@ IV-C Evaluation Results
In the weather dataset, the SVM operator results from sub-figures 5(a) and 5(b) show that using $300$ vectors in the representation space consistently led to more accurate operator models across all similarity methods. Specifically, cosine similarity in combination with the $300$-dimensional vector embedding reduced the error rate in operator predictions, demonstrating that projecting datasets into this representation space and applying cosine similarity improves the prediction accuracy on the modelled operator. For the MLP classifier (Figures 5(c) and 5(d)), the results illustrate that using vector embeddings of size $300$ and Cosine similarity- produced the most accurate MLP classifier operators.
(a) SVM with SGD RMSE error loss
(b) SVM with SGD MAE error loss
(c) MLP RMSE error loss
(d) MLP MAE error loss
@@
(a) SVM with SGD RMSE error loss
(b) SVM with SGD MAE error loss
(c) MLP RMSE error loss
(d) MLP MAE error loss
Figure 6: Adult dataset prediction error loss
@@
TABLE II: Evaluation results of our framework exported analytic operator with lowest prediction error in comparison with Apollo
@@
Table II compares model operators, loss functions, and speedup metrics for our framework and Apollo at SR values of $0.1$ and $0.2$. Methods $100$V, $200$V, and $300$V denote vector embedding dimensions. The lowest prediction errors align with our pipeline’s similarity search method. Apollo outperforms our framework on the Stock dataset for the LR analytic operator at SR equals with $0.1$ (with $50\%$ and $6\%$ improvements for RMSE and MAE, respectively). However, our framework excels with the MLP regression operator, improving RMSE and MAE by $20\%$ and $17\%$, respectively. The LR operator’s performance gap on the Stock dataset is minor. For other datasets, our framework consistently surpasses Apollo across different SR values. This demonstrates the effectiveness of our similarity search approach, which enhances data quality and reduces $\Phi$ prediction errors by identifying relevant datasets $D_{r}$ from the data lake directory $D$. The Adult dataset also highlights our framework’s advantage with increasing feature dimensions. Although Apollo achieves better raw speedup due to the higher complexity of our framework’s vectorization step, our framework outperforms it in amortized speedup. By excluding the reusable vectorization process, it achieves speed gains of $10\%$ to $60\%$ for most operators. The SR approach, leveraging vector embedding representations, enhances operator prediction compared to Apollo and achieves greater amortized speedup. However, the similarity search method outperforms both Apollo and the SR approach in prediction accuracy and amortized speedup, establishing its clear superiority across most datasets and operator scenarios.
IV-D NumTabData2Vec Evaluation Results
@@
TABLE III: Similarity between vectors of different datasets scenarios
Model Name Similarity
NumTabData2Vec $100$ Vector size $0.54$
NumTabData2Vec $200$ Vector size $0.18$
NumTabData2Vec $300$ Vector size $0.16$
Our proposed model, NumTabData2Vec, was evaluated to determine its ability to distinguish dataset scenarios based on qualitative differences. For each scenario, $n$ random datasets were selected, and their vector embeddings averaged to represent the scenario. These embeddings, initially in a 200-dimensional space, were projected into 3D using PCA and are shown in Figure 8. The figure illustrates that each dataset scenario occupies a distinct space, with minimal overlap or clustering. This demonstrates that NumTabData2Vec effectively distinguishes datasets, outperforming prior methods like Word2Vec and Graph2Vec [14, 15], which achieved similar accuracy but on different data types (e.g., words, graphs) rather than entire datasets. Table III further highlights the average cosine similarity between dataset embeddings, showing greater dissimilarity as vector dimensions increase. However, results suggest that dimensions between $100$ and $300$ are sufficient for accurate distinction, avoiding the need for larger vector sizes.
@@
TABLE IV: NumTabData2Vec execution time for different dataset dimensions and different vector sizes
To evaluate how dataset dimensions affect the execution time of NumTabData2Vec, we created synthetic datasets with varying numbers of rows ($100$, $500$, $1000$, $1500$, and $1800$) and columns ($3$, $10$, and $20$). These datasets were vectorized into different dimensions, and the execution times were recorded. Table IV shows that increasing the k-dimension requires approximately $20\%$ more time to generate the vector embeddings. This is expected, as a higher k-dimension involves more hyperparameters, which naturally increases computation time.
@@
To evaluate NumTabData2Vec’s ability to distinguish datasets based on different properties like distribution and order, we introduced Gaussian noise to random $l\%$ of data tuples in an HPC dataset. Figure 10 visualises the original and noise-modified datasets, projected from a 200-dimensional space to 2D using PCA. Each bullet caption g denotes the percentage of Gaussian noise added in the dataset. As noise increases, the representation space shifts further from the original dataset, indicating that NumTabData2Vec effectively captures distribution differences. Additionally, since the HPC dataset has an inherent order, the model’s sensitivity to noise demonstrates its ability to distinguish datasets based on ordering as well.
To evaluate how fine-grained as distinction can be, we introduced noise into a single column and repeated the previous experiment, with the difference being that noise was added exclusively to the first column. Figure 11 visualizes the dataset’s 2D vector space. The amount of Gaussian noise added to the dataset’s first column is indicated by g in the bullet caption. The results show that as more noise is introduced to the column, the vector representation moves further away from the original dataset. In contrast to the previous experiment shown in Figure 10, the noisy dataset’s representation stays closest to the original when only a single column is modified. Also in this experiment the dataset points in the 2-dimension are more grouped between them instead the previous experiment.
V Limitations
There exist a number of limitations in our work as we described it. In this section we briefly highlight them. Firstly, our input datasets comprise records of specific size and type (numerical). This currently excludes data with textual and categorical attributes, or tables with varying number of features inside a set of datasets. Secondly, we currently consider single-input and single-output operator modelling. Finally, our proposed NumData2Vec model for data vectorization has a performance limitation, as it cannot deal with datasets bigger than about $3000$ tuples. This is mostly a hardware limitation of off-the-shelf GPUs (with at most 24GB of memory available).
VI Conclusion

### v3

@@ 3.3 Dataset Selection
The selection of the most similar datasets has been implemented using three different approaches. Different similarity functions are easily plugged into our pipeline. The first method uses cosine similarity, which measures the angle between two vectors in the embedding space $z$ independent of their magnitudes, with a higher value indicating greater similarity. The alternative method calculates the Euclidean distance between two vectors to capture their geometric closeness and determine dataset similarity. This approach aids in selecting the most relevant datasets according to organizational requirements. The smaller the distance value then more similar are the datasets. Dataset selection using cosine similarity or euclidean distance selects a fraction of $\lambda,\lambda>0$ of the closest datasets to dataset $D_{o}$. The third approach involves utilising the K-Means [19, 21] clustering technique to choose the most relevant datasets. The datasets from $D$ are divided into $s$ ($s>1$) separate clusters, where datasets with similar features are grouped together in the same cluster based on similarity equations. We determine the optimal number of clusters using silhouette scores [27]. This is done by the following equation:
$SilhouetteScore(z_{i})=\frac{b(z_{i})-a(z_{i})}{\max(a(z_{i}),b(z_{i}))},\$ (5)
@@
where for each vector $z_{i}$ computes the mean intra-cluster distance ($a(z_{i})$) which is the distance with the other vectors in the same cluster, and the mean nearest-cluster distance ($b(z_{i})$) is the minimum average distance with the other vectors in a different cluster. Silhouette Score ranges from $-1$ to $1$, and the higher value defines the best $s$ number for clusters.
@@
Algorithm 1, outlines the K-means process for selecting relevant datasets $D_{r}$ based on the target dataset $D_{o}$. Using the optimal s (with the highest Silhouette score), the vector representations $z$ of each dataset are clustered. Next, the algorithm uses the vector $z_{o}$ of dataset $D_{o}$ to find the closest cluster centroid. All datasets in that cluster are defined as the relevant datasets $D_{r}$, which are then used to model the analytics operator. However, we defined a maximum and minimum size for $D_{r}$, and if these conditions are not met, datasets are either removed from the cluster or added based on their distance from the cluster centroid. These small adjustments are only made in cases where the clustering technique does not yield results that satisfy our requirements.
4 Evaluation
@@
We compared our framework with Apollo [9, 4], which models analytic operators using data content. Two loss functions to measure prediction accuracy are employed: root-mean-square error (RMSE) and mean absolute error (MAE). RMSE is sensitive to outliers, while MAE is not; conversely, RMSE accounts for error direction, which MAE cannot. We further assess efficiency using Speedup and Amortized Speedup metrics, where Speedup is defined as $\frac{T{{}^{(i)}_{op}}}{T{{}^{(i)}_{SimOp}+T_{vec}+T_{sim}+T_{pred}}}$, where $T{{}^{(i)}_{op}}$ is the time to execute operator $i$ on all datasets, $T{{}^{(i)}_{SimOp}}$ is the time to model the operator with datasets from similarity search, $T_{vec}$ is the vector embedding computation time, $T_{sim}$ is the similarity search time, and $T_{pred}$ is the prediction time for $D_{o}$. Amortized speedup including one-time vectorization per data lake across multiple operators. Three variants with vector sizes 100, 200, and 300 (each with eight transformer layers) were trained for $100$ epochs on four NVIDIA A10101010 GPUs. More experimental evaluation results can be found in the extended version of this work [20].
4.1 Evaluation Setup
@@
Our framework is deployed over an AWS EC2 virtual machine server running with 48 vCPUs of AMD EPYC 7R32 processors at 2.40GHz, and four A10s GPUs with 24GB of memory each, $192GB$ of RAM memory, and $2TB$ of storage, running over Ubuntu 24.4 LTS. Our code is written in Python (v.3.9.1) and PyTorch modules (v.2.4.0). Apollo was deployed on the same AWS EC2 virtual machine server, utilizing only the vCPUs and RAM, as it does not require a GPU for execution.
4.2 Datasets
Table 1: Dataset properties for experimental evaluation
@@ 4.2 Datasets
We evaluated our framework using four real-world datasets (see Table 1). The NumTabData2Vec module was trained on separate data ($60\%$ training, $40\%$ testing). The Household Power Consumption (HPC) dataset [13] contains $401$ datasets with $2051$ tuples and seven features recorded at one-minute intervals of electric power usage measurements. The Adult dataset [5], used for binary classification, predicts income levels and includes $100$ datasets with $228$ individuals and socio-economic features. The Stock Market dataset [1] consists of $508$ datasets with $13$ to $1959$ tuples describing daily NASDAQ stock prices. Weather dataset [2] provides hourly measurements from $36$ U.S. cities (2012–2017), split into $49$ datasets with $516$ tuples and seven features. Any categorical feature column in all datasets is transformed to numerical data by one-hot encoding. These datasets were selected to demonstrate our framework’s ability to perform consistently across diverse real-world scenarios.
Our framework was evaluated by predicting the outputs of various ML operators without directly executing them. Datasets were projected into $k$-dimensional spaces with vector dimensions of $100$, $200$, and $300$. For regression datasets (Household Power Consumption and Stock Market), we modelled Linear Regression (LR) and Multi-Layer Perceptron (MLP), while for classification datasets (Weather and Adult), we modelled Support Vector Machine (SVM) and MLP classifiers. Each experiment has executed 10 times, and we report the average error loss and speedup.
4.3 Evaluation Results
Figures 4, 4, 6, and 6 present the evaluation results of different similarity search methods across vector embedding spaces of sizes $100$, $200$, and $300$ (green, blue, and grey bars, respectively). In each sub-figure, the y-axis represents the error loss value, while the x-axis displays the similarity search method applied over the vector embeddings. Figures 4 and 4 display results for the Stock Market and Household Power Consumption datasets, with MLP regression in the bottom sub-figure and LR in the top. Figures 6 and 6 show results for the Weather and Adult datasets, with SVM (SGD) in the top sub-figure and MLP classifier in the bottom. Left sub-figures use RMSE loss, while right sub-figures use MAE loss.
@@ 4.3 Evaluation Results
Figure 4, for the HPC dataset, shows as increase the vector dimension size there is slightly lower prediction error for all the operator modelling. Different similarity methods do not result in any significant differences in the prediction error loss for all the operator modelling. This suggests that, regardless the similarity selection method, our framework effectively selects the most optimal subset of data to improve model predictions on the unseen input dataset $D_{o}$. Additionally, we observe higher error loss with a vector size of 100, which can be attributed to the reduced representation capacity of lower-dimensional vectors. This limitation results in fewer “right” datasets being selected.
@@
In the weather dataset, the SVM operator results (Figures 5(a) and 5(b)) show that using $300$ sized vectors in the representation space consistently led to more accurate operator models across all similarity methods. Specifically, cosine similarity in combination with the $300$-dimensional vector embedding reduced the error rate in operator predictions, demonstrating that projecting datasets into this representation space and applying cosine similarity improves the prediction accuracy on the modelled operator. For the MLP classifier (Figures 5(c) and 5(d)), the results illustrate that using vector embeddings of size $300$ and Cosine similarity- produced the most accurate MLP classifier operators.
@@
Figure 6: Adult dataset prediction error loss
@@
Table 2: Evaluation results of our framework exported analytic operator with lowest prediction error in comparison with Apollo
@@
For the HPC and Weather datasets, the SR approach was approximately $15\%$ less accurate in operator prediction compared to all similarity search methods, even as vector embedding dimensions increased. In contrast, the Stock dataset exhibited a significantly larger discrepancy, with the SR approach performing about $70\%$ worse in prediction accuracy across all vector embedding dimensions. Similarly, in the Adult dataset, the SR approach recorded the poorest performance, delivering nearly $90\%$ worse prediction accuracy compared to the similarity search methods.
Table 2 compares model operators, loss functions, and speedup metrics for our framework and Apollo at SR values of $0.1$ and $0.2$. Methods $100$V, $200$V, and $300$V denote vector embedding dimensions. The lowest prediction errors align with our pipeline’s similarity search method. Apollo outperforms our framework on the Stock dataset for the LR analytic operator with the smallest amount of SR. However, our framework excels with the MLP regression operator, improving RMSE and MAE by $20\%$ and $17\%$, respectively. The LR operator’s performance gap on the Stock dataset is minor. For other datasets, our framework consistently surpasses Apollo across different SR values. This demonstrates the effectiveness of our similarity search approach, which enhances data quality and reduces $\Phi$ prediction errors by identifying relevant datasets $D_{r}$ from the data lake directory $D$. The Adult dataset also highlights our framework’s advantage with increasing feature dimensions. Although Apollo achieves better raw speedup due to the higher complexity of our framework’s vectorization step, our framework outperforms it in amortized speedup. By excluding the reusable vectorization process, it achieves speed gains of $10\%$ to $60\%$ for most operators. The SR approach, leveraging vector embedding representations, enhances operator prediction compared to Apollo and achieves greater amortized speedup. However, the similarity search method outperforms both Apollo and the SR approach in prediction accuracy and amortized speedup, establishing its clear superiority across most datasets and operator scenarios.
4.4 NumTabData2Vec Evaluation Results
Table 3: Similarity between vectors of different datasets scenarios
NumTabData2Vec Vector Size Similarity
$100$ $0.54$
$200$ $0.18$
$300$ $0.16$
@@ 4.4 NumTabData2Vec Evaluation Results
Our proposed model, NumTabData2Vec,was evaluated for its ability to distinguish dataset scenarios based on qualitative differences. For each scenario, $n$ random datasets were reduced to 3D vector embeddings using PCA, as shown in Figure 9, which demonstrates NumTabData2Vec’s ability to distinguish datasets with minimal overlap across contexts. Unlike prior methods [22, 24], which focus on text or graphs, NumTabData2Vec applies to entire datasets. Table 3 further highlights the average cosine similarity between dataset embeddings, showing greater dissimilarity as vector dimensions increase. However, results suggest that dimensions between $100$ and $300$ are sufficient for accurate distinction, avoiding the need for larger vector sizes.
Table 4: NumTabData2Vec execution time for different dataset dimensions and different vector sizes
@@
To evaluate NumTabData2Vec’s ability to distinguish datasets by distribution and order, Gaussian noise was added to $l\%$ of tuples in an HPC dataset. Figure 11 shows the original and noise-modified datasets projected to 2D using PCA, with greater noise causing larger shifts in the representation space. This demonstrates the model’s effectiveness in capturing distribution differences and distinguishing datasets based on ordering.
To assess fine-grained distinctions, we repeated the experiment by adding Gaussian noise exclusively to the first column of the dataset. Figure 11 shows the 2D vector space, where g in the bullet caption indicates the noise level. As noise increases, the representation shifts further from the original dataset, though it remains closer than in Figure 11, with points more tightly grouped in 2D space.
To evaluate how dataset dimensions affect the execution time of NumTabData2Vec, we created synthetic datasets with varying numbers of rows ($100$, $500$, $1000$, $1500$, and $1800$) and columns ($3$, $10$, and $20$). These datasets were vectorized into different dimensions, and the execution times were recorded. Table 4 shows that increasing the k-dimension requires approximately $20\%$ more time to generate the vector embeddings. This is expected, as a higher k-dimension involves more hyperparameters, which naturally increases computation time.
@@
5 Limitations
There exist a number of limitations in our work as we described it. In this section we briefly highlight them. Firstly, our input datasets comprise records of specific size and type (numerical). This currently excludes data with textual and categorical attributes, or tables with varying number of features inside a set of datasets. Secondly, we currently consider single-input and single-output operator modelling. Finally, our proposed NumData2Vec model for data vectorization has a performance limitation, as it cannot deal with datasets bigger than about $3000$ tuples. This is mostly a hardware limitation of off-the-shelf GPUs (with at most 24GB of memory available for a budget GPU).
6 Conclusion